OpenStudy (anonymous):

Need help please, thanks (: Given z = 8 and x = 6, what is the ratio of z to x?

OpenStudy (anonymous):

4/3

OpenStudy (anonymous):

z:x z/x=8/6=4/3

OpenStudy (anonymous):

thanks y'all

OpenStudy (anonymous):

anytime